OpenStudy (anonymous):

Give an intuition as to why 1/2 + 1/4 + 1/8 + 1/16 + ..... = 1

OpenStudy (amistre64):

becasue you are slowly adding smaller values

OpenStudy (amistre64):

if I start from here and work myself half way each time; i will eventually get close to you, even tho I may never reach you

OpenStudy (anonymous):

I don't think that's enough, the series 1/2 + 1/3 + 1/5 ... doesn't add to 1.

OpenStudy (amistre64):

lol ... you asked for intuition :)

OpenStudy (anonymous):

Yes, but I meant something that would be accurate with the series converging to the value of 1. But I get your point.

OpenStudy (anonymous):

Maybe think of a square that is 1 by 1. Now color in 1/2 the area. You have the other 1/2 blank. Now color in 1/2 of the remaining (which is 1/4). Continue to color in half of what is remaining and you'll get closer and closer to coloring in the entire square, which has area = 1.

OpenStudy (anonymous):

amistre is absolutely right... if u see 1/2 is half of one, so u hv covered 1/2 the distance to 1 then 1/2 of 1/2 = 1/4 , so u hv covered 3/4 of total distance to 1 then 1/2 of 1/4 = 1/8 and so on u will get very very very near to one but never to one theoretically however for practical purposes u will hv reached 1

OpenStudy (anonymous):

joebrown: That's what I was looking for. :) But everybody gets a medal.

OpenStudy (amistre64):

oh ... a geometric interpretation ;) my confusion was in the word "intuition" :)

OpenStudy (anonymous):

Now go to my other question. Of why 0. 999999.... and 1 are the same.

OpenStudy (anonymous):

whatever route one take, this will also lead to the square being "nearly" filled and not completely filled.....

OpenStudy (amistre64):

if i take a carrot and chop it in half each time, I will get ever so close to chopping up the whole thng

OpenStudy (amistre64):

there are a few reasons why .9999.... equals 1

OpenStudy (amistre64):

the obvious is that 1/3 is a real number that decimates to .3333.......

OpenStudy (amistre64):

1/3 + 1/3 + 1/3 = 1 .333... + .333... + .333... = .999... = 1

OpenStudy (anonymous):

If you let x = 0.9999999... Then that means that 10x = 9.999999.... Then 10x - x = 9.99999... - 0.9999... And 9x = 9 So x = 1.

OpenStudy (anonymous):

Those are both great ways of proving it. ^^

OpenStudy (anonymous):

as they say "A cat can be skinned in more than one way....."

OpenStudy (amistre64):

.99999 = 9(.1) + 9(.01) + 9(.001) + ...

OpenStudy (amistre64):

\[\frac{ar}{1-r}=1\] \[\frac{9(1/0)}{1-(1/10)}=1\]

OpenStudy (amistre64):

ack .... typing and mathing dont mix

OpenStudy (anonymous):

We really had some brain storming here. :D